OpenStudy (anonymous):

You deposit $4,000 at the end of each year into an ordinary annuity with an interest rate of 5% compounded annually. What will be the value of the annuity after 10 years? Do not round until the final answer; then round to the nearest dollar as needed.

OpenStudy (wolf1728):

The formula for calculating an annuity is: Total = Yrly Amt *([(1+r)^(n+1) -1] / r) - Yrly Amt Total = $4,000.00 * ([(1+.05)^11]-1 / .05) - $4,000.00 Total = $4,000.00 * (.7103393581 / .05) - $4,000.00 Total = ($4,000.00 * 14.206787162) - 4,000.00 Total = 56,827.15 - 4,000.00 Total = 52,827.15 Calculator to check the work: http://www.1728.org/annuity.htm

OpenStudy (wolf1728):

Oh and rounded to the nearest dollar it is: $52,827